What are the symmetry operations possible for the following structures, and what is the expected point group?

a). The tetrahedral CoCl42-

b). Cl-I-Cl-

c). C6H12 (chair form)  
Answer #1

Part A

CoCl_4^{2-} is a symmetric tetrahedral molecule. It belongs to a special point group T_d. The symetry operations of the point group are,

Identity (E)

Three fold rotation axis (C_3)

Two fold rotation axis (C_2)

Four fold improper rotation axis (S_4)

Dihedral mirror planes (\sigma _d)

Part B

ICl_2^- is a linear molecule. This belongs to a special point group, D_{\infty h}. The operations possible are,

Infintite rotational axes of any fold (C_\infty)

Centre of inversion (i)

Infintite improper rotational axes of any fold (S_\infty)

An infininite number of two fold rotational axes through the centre of inversion (C_2)

An infinite number of verticle mirror planes (\sigma _v)

Part C

The chair conformation of cyclohexane belongs to the D_{3d} point group and has the following operations.

Identity (E)

Three fold rotation axis (C_3)

Two fold rotation axis (perpendicular to the 3-fold axis) (C_2)

Centre of inversion (i)

Six fold improper rotation axis (S_6)

Dihedral mirror planes (\sigma _d)